The processors of Peanut's Butter indicate on the label that the bottle contains 16 ounces of peanuts. The standard deviation of the process is 0.5 ounces. A sample of 36 bottles from the last hour's production revealed a mean weight of 16.12 ounces per bottles, at the α = 0.05 is the process out of control? That is, can we conclude that the mean amount per bottle is different from 16 ounces. Find the critical value?


1
Expert's answer
2021-05-25T10:02:03-0400

We have given that,

"\mu = 16"

"\sigma = 0.5"

"n = 36"

"\bar x = 16.12"

"\alpha= 0.05"

"H_0: \mu_0 = 16"

"H_0 : \mu_0 \ne 16"

Applying z test,

"z = \dfrac{16.12-16}{\dfrac{0.5}{\sqrt{36}}}"


"z = \dfrac{0.12}{0.083}" ; "z = 1.44"


We know that value of "z" at 0.05 level of significance the critical value "=\pm 1.96"

Hence, our calculated value is less than te absolute value of "z_{\alpha}" .

Therefore, Null Hypothesis is verified.